OpenStudy (anonymous):

The value of y varies directly with x, and y = - 18 when x = 6. Find y when x = - 8.

OpenStudy (mertsj):

y=kx is the equation for a direct variation. Plug in y and x and find k.

OpenStudy (mertsj):

What did you get for k?

OpenStudy (mertsj):

It's very difficult to help you if you won't answer me.

Directrix (directrix):

y=kx -18 = k(6) -3 = k y = (-3)* (-8) y = 24

OpenStudy (anonymous):

Ohh dang my bad, was watching a movie